# kGym: A Platform and Dataset to Benchmark Large Language Models on Linux Kernel Crash Resolution

### v3

@@ 7 Conclusion
In this work, we introduce kBench, a new challenging SE benchmark aimed at Linux kernel crash resolution. To effectively experiment on kBench, we also introduce kGym, a platform to execute large-scale kernel experiments. To interact with kGym, we also provide few simple APIs. Using kGym, we run over $17$k kernel jobs to report our initial baseline results that indicate poor performance even when using state-of-the-art LLMs. Thus, we conclude that there is adequate scope for research to improve crash resolution performance in massive production-ready systems-level codebases. We hope that by introducing kBench and kGym, we spur more efforts that lower the barrier of entry to research at the intersection of machine learning and system software.
@@ A.3 Prompt Template
A.4 Subset of kBench for every model
Table 10: For each LLM, the final subset of bugs from the kBench depends on the chosen retrieval method and the maximum allowed context length.

### v4

@@ 8 Conclusion
In this work, we introduce kBenchSyz (and kBenchC), a new challenging SE benchmark aimed at Linux kernel crash resolution. To effectively experiment on kBenchSyz, we also introduce kGym, a platform to execute large-scale kernel experiments. To interact with kGym, we also provide few simple APIs. Using kGym, we run over $17$k kernel jobs to report our initial baseline results that indicate poor performance even when using state-of-the-art LLMs. Thus, we conclude that there is adequate scope for research to improve crash resolution performance in massive production-ready systems-level codebases. We hope that by introducing kBenchSyz and kGym, we spur more efforts that lower the barrier of entry to research at the intersection of machine learning and system software.
@@ A.3 Prompt Template
A.4 Subset of kBenchSyz for every model
Table 10: For each LLM, the final subset of bugs from the kBenchSyz depends on the chosen retrieval method and the maximum allowed context length.